OpenStudy (anonymous):

A cone is inscribed in a regular square pyramid. If the pyramid has a base edge of 6" and a slant height of 9", find the volume of the cone. 12√2 18√2 27

OpenStudy (anonymous):

hey i think your answer is this your answer is 25.47 in^3

OpenStudy (anonymous):

area of the cone will equal 1/3 (area of the base) (height) for the area of the base, we know that the cone has a diameter of 6" (since it is the largest possible circle inside the square base). so use the area of a circle formula to get: (pi)r^2 where r=3 (1/2 of 6) 9(pi) the height is a little tickier, since you know the slant height, you need to use the properties of a right triangle to get the height. The hypotenuse is 9" and the base is 3" so we use a^2 + b^2 = c^2 3^2 + b^2= 9^2 9+b^2=81 b^2=72 b=8.49 now that you have the height and the area of the base, put it all together to get: 1/3 (9(pi))(8.49) your answer is 25.47 in^3

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

Thanks, again(:

OpenStudy (anonymous):

umm no thank you(:::